# Show Me What and Where has Changed? Question Answering and Grounding for Remote Sensing Change Detection

### v1

@@ 6 Conclusions and Limitations
In this paper, we addressed key limitations of the classic CDVQA task, i.e., its inability to simultaneously provide textual answers and corresponding visual evidence. To overcome this, we proposed the CDQAG benchmark, designed to offer intuitive visual explanations alongside textual answers. To facilitate research on CDQAG, we constructed the first large-scale dataset QAG-360K, and proposed a effective baseline method VisTA, which achieves the new SoTA performance on both our QAG-360K and CDVQA datasets. We hope that our work can shed new light on change detection. However, this study has two limitations: (1) The limited availability of public change detection datasets restricts the growth of this field. (2) While MLLMs show potential in complex reasoning, our work only establishes a baseline without an in-depth exploration of their performance. Future work could explore advanced MLLM architectures for CDQAG, which may yield deeper insights and improved results.

### v2

@@ 6 Conclusions and Limitations
In this paper, we proposed the CDQAG benchmark, designed to provide intuitive visual explanations alongside textual answers. To facilitate research on CDQAG, we developed a large-scale dataset QAG-360K, and introduced a powerful baseline method VisTA, which sets new state-of-the-art performance on both the CDVQA and our QAG-360K datasets. In addition, we present several analyses and insights for future research: (1) Our experiments demonstrates that effective multimodal interaction is essential for analyzing diverse scenarios in change detection. This necessitates more efforts in achieving reliable complex reasoning. (2) While our solution enhances question answering and grounding, delivering precise visual answers persists as a formidable challenge. Exploring solutions to this will greatly benefit the CDQAG task. Finally, we highlight the importance of CDQAG for practical applications in remote sensing and hope it can inspire the community to drive progress in this domain.